Use inductive reasoning to predict the next three numbers in the pattern:
a. 0,1,1,2,3,5,8___,____,_____
b.-12,-7,-2,3,8,____,______,_____
c. 1/5,1/3,3/7,1/2,5/9,_____,______,_______

Answer by edjones(8007)

a)13,21,34 take the sum of the 2 preceding numbers.
b)13,18,23 Add 5 to each number.
c)6/10=3/5,7/11,8/12=2/3 Add 1 to both the numerator and denominator each time.
.
Ed